Fill in the blank

The number of possible hosts in the subnet whose subnet mask is 136.76.29.0/22 is ______ . _________________ is the mechanism where an IPv6 packet is sent over an IPv4 network of routers and links. . Most of the link layer is implemented in a ________________________ . ________________________ refers to the ability of the receiver to both detect and correct errors. . The IP address to MAC address translation is provided by ________________________ . The MAC address is a ______________ layer address. . The ______________________ does the translation between IP addresses and MAC addresses. . ________________ is the most commonly used wired LAN network. The packet processing is done only at the Link and Physical layers in a _____________ . The network interface card is also known as _________________ . The sizes of the address spaces of IPv4, IPv6, and MAC address are (respectively) ________, ________, and ________ . The ability of the receiver to both detect and correct errors is known as _____________________
Answer #1

Dear Student,

Here are the answers....

---------------------------------------------------------------------------------------------------------------------------------------

The number of possible hosts in the subnet whose subnet mask is 136.76.29.0/22 is = 210 = 1024 -2 = 1022 hosts. (because 22 bits are reserved for network address and 10 bits for host address.).

6to4 is an Internet transition mechanism where an IPv6 packet is sent over an IPv4 network of routers and links.

Forward Error Correction (FEC) refers to the ability of the receiver to both detect and correct errors.

The IP address to MAC address translation is provided by ARP(Address Resolution Protocol) Protocol.

The MAC address is a Physical layer address.

The ARP(Address Resolution Protocol) protocol does the translation between IP addresses and MAC addresses.

Ethernet is the most commonly used wired LAN network.

The packet processing is done only at the Link and Physical layers in a OSI 7 layer Model.

The network interface card is also known as Network Interface Controller.

The sizes of the address spaces of IPv4, IPv6, and MAC address are (respectively) 32 bytes, 128 bytes, and 48 bytes.

The ability of the receiver to both detect and correct errors is known as Forward Error Correction (FEC).

------------------------------------------------------------------------------------------------------------------------------------

Kindly Check and Verify Thanks...!!!